### v1

# Speed-Accuracy Trade-Off in the Simulation of Realistic Arterial Blood Flow: Effects of Geometric Simplifications and Non-Newtonian Models

Abstract
This study investigates various aspects of blood flow in realistic arteries obtained from Computed Tomography (CT) scan data and its geometrically simplified variants. The study uses realistic inflow and outlet boundary conditions with Newtonian and non-Newtonian blood viscosity models, including the one recently developed by Apostolidis and Beris [28]. Similar studies exist to predict the progression of atherosclerosis, clots, and thrombus formation in arteries. However, these assume smooth flow geometries. In contrast, realistic arteries possess significant amounts of roughness, perturbations, and undulations. To understand their impact, studies are performed with two levels of geometry simplifications. Our results indicate that flow field predictions on these simplified geometries are visibly affected. The first level of simplification loses the local asymmetry but is able to reasonably approximate various parameters, capturing patterns with the correct magnitude. However, further simplification to a smooth one does not show vortex structures in the flow field. The flow recirculation zones near wall perturbations progressively weaken as the geometry is smoothed. On the other hand, no significant differences are observed between Newtonian and non-Newtonian models in this study. Our analysis also enables a speed-accuracy trade-off for future applications. The number of iterations to convergence becomes progressively lower as geometry is simplified, with nearly four times the difference obtained even with the first level of simplification. Further, the computational time for the Newtonian model is nearly 10 times smaller than the non-Newtonian counterpart.
Keywords— blood flow, descending aorta, geometry perturbations, pulsatile waveform, wall shear stress
1 Introduction
Cardiovascular diseases (CVD) remain one primary reason for human mortality [1] and have grown in importance after the COVID-19 pandemic. The progression of CVD is directly related to the geometric details of the arteries [2]. CVD is also known to aggravate other conditions as well [3]. Over the last few decades, greater computing power has enabled detailed computational studies on blood flow in the human arteries [4, 5]. Researchers have been interested in the link between atherosclerosis and fluid flow and related parameters, including the wall shear stress (WSS) [6, 7, 8, 9]. Increasingly, studies have used patient-specific geometries with the aim of contributing to medical applications [10]. Similar studies have been reported to predict the progression of atherosclerosis [11, 12], clots [13], and formation of thrombus [14]. Significant studies have investigated the effect of stenosis on blood flow [15, 16, 17]. However, in all these studies, the flow geometries have been assumed to be quite regular. In contrast, human arteries are far from regular and possess a significant amount of undulations in diameter as well as in the flow direction. The geometrical variations of arteries have been linked to the onset of atherosclerosis [18, 19, 20], thereby justifying its importance in the broader aspect of human health.
Despite the apparent importance of irregularities in geometry, researchers have rarely focused on their blood flow. Yi et al. [21] performed simulations to study the impact of surface roughness on arterial stenosis. Shukla et al.[22] studied the effect of surface roughness on pulsatile flow in a tube with a periodic pattern of roughness in the form of sinusoids while Owen et al. [23] studied the effect of surface roughness on blood flow in a tube. To the best of our knowledge, these three studies come close to studying the impact of roughness from a biological perspective for blood flow analysis. Among them, the studies by Yi et al. [21] and Shukla et al. [22] use Newtonian fluid models. Owen et al. [23] used a variety of models, including Newtonian and non-Newtonian ones, to study the effect of an uneven geometry. However, the geometry and setup they created in the study can be argued to be far removed from reality. In a tube, they made a ”rough” patch with surface irregularities that extended along the length and compared fluid flow results between the rough patch and an equivalent smooth patch on the other side. This geometrical setup has two significant problems. Firstly, the roughness is distributed all over the surface, and a tube with a rough and smooth patch is not close to the surface features of arteries in our body. Secondly, the surface features comprise not just roughness but also changes on the scale of the artery diameter. Hence, irregularities also exist in the shape and size fluctuations along the axial length of the arterial tube. The flowing fluid experiences the combined effect of these irregularities. As discussed later, the real artery may deviate significantly from the circular shape over a given plane. Additionally, the equivalent diameter can also show variations along the axis. To the best of our knowledge, no previous study has comprehensively investigated the effect of such shape and size perturbations, in addition to surface roughness, which is aimed for in the present study. Also, we use a variety of viscosity models to examine the sensitivity of the non-Newtonian nature of blood with geometric perturbations.
Thus, in this study, we start with a patient-specific geometry – a part of the descending aorta with all possible details and perturbations retained for blood flow calculations. To mimic real-life conditions, we use a variety of viscosity models and realistic pulsatile inflow and pressure outlet boundary conditions. Recent developments in computational fluid dynamics have enabled blood flow studies in realistic arterial geometries. Thus, it is possible to compute flow distribution in three dimensions for pulsatile flow in branches and bifurcations with bulges and constrictions. When the geometry is made realistic, undulations and perturbations of the artery wall need refinement in mesh size, thereby increasing computational cost. Secondly, for an accurate representation of a complex fluid such as blood, non-Newtonian rheology models need to be used, further increasing complexity. In the present study, we explore the consequences of simplifying the geometry in the context of various viscosity models (with an additional aim of reducing computational cost). The original artery model is obtained from CT scans, and progressively simpler geometries are generated by smoothing operations in image processing. We study hemodynamic aspects using Newtonian and non-Newtonian viscosity models in these equivalent geometries. Further, comparisons of results across these different geometries can yield insights into the importance of geometric resolution and perturbations on the flow field.
Towards this goal, three levels of equivalent geometries are studied. Level 1 is the original geometry derived from DICOM images provided by a medical hospital. Level 2 represents the same artery, divided into slices along its length, each slice representing a smooth duct of constant diameter. The number of slices can be varied to create varying levels of overall “smoothness”. Level 3 is a straight, smooth, circular tube of linearly varying diameter whose inlet and outlet diameters match the original. Inlet flow waveform is specified in terms of the peak flow rate and frequency. The significance of the outcome of this study is manifold. The most direct one is linked to medical measurements and device points of view. Alternatively, the study provides us with an understanding of the modeling aspects required for a given level of resolution of the selected artery. If we need to perform CFD calculations on an arterial geometry, the question of interest is, will a simplified geometry be sufficient? Or will the actual geometry with all small-scale details be needed to obtain medically relevant results? In addition, such studies also reveal the speed-accuracy trade-off. It is expected that the simplified geometric representation can be simulated in a lesser amount of computational time. However, some information is bound to be lost in the process of geometric simplification. For a particular application, one needs to anticipate the compromise in geometry permitted to perform a sufficiently accurate calculation in as little time as possible. This question is addressed in the present study by adopting equivalent geometries as described earlier.
One additional aspect studied here is the change obtained by the choice of the viscosity model. Calculations in these three equivalent geometric representations provide us with intricate details of the effects of geometric perturbations on the flow of blood and their connection with the chosen rheological model, apart from the effect of these details on the computational speed.
The outcomes of the simulations are shown in terms of vortical patterns and hemodynamic parameters such as the time average wall shear stress (TAWSS) and oscillatory shear index (OSI). The differences in flow patterns among the three levels are discussed in detail, with an understanding of the reduction in CPU time, which is substantial. The results presented in this study may enable the simplification of complex arterial networks to increase the feasibility of meaningful full-body simulations in real-time.
2 Numerical Methodology
2.1 Geometry
The actual geometry of a descending aorta considered here is from patient-specific CT scan data obtained from the LPS Institute of Cardiology, Kanpur, India. The patient studied is a 40-year-old female and is otherwise anonymous. The reconstruction of the descending aorta is performed by the open-source software SimVascular v-2017 [24], in which the pathline of the center of the artery is traced, followed by segmentation of the peripheral region of the artery and surface generation. The selected artery length and details of the cross-sectional areas are given in Table 1. For the present study, the descending aorta is simplified into three levels. Level 1 is the original patuent-specific artery created by the DICOM images arising from a CT scan. For the second level (simplified relative to the original, which we denote as Level 2), the original artery, i.e., Level 1, is divided into 25 slices, and a smooth CAD geometry is constructed based on the average diameter of the slices and is shown in Fig 1.
Figure 1: Level of geometrical aspects is shown, in (i) Level 1 of descending aorta extracted from CT-scan data, (i) Slice-based geometry of descending aorta (Level 1 is divided into 25 slices for the construction of level 2). (iii) Level 3 geometry based on the inlet and outlet diameter of Level 1 geometry. (iv) Shows the cross-section area variation over the normalized z-axis for each level of descending aorta simplification.
Level 2 is a simplified geometry that contains features of the original but is much coarser with fewer details. Next, we consider a highly simplified version that is a tapered smooth tube based on the inlet and outlet diameter, which we denote as Level 3. The three levels considered here represent a hierarchical simplification of the original geometry, retaining partial features of the original. It can be hypothesized that the grid requirements (and thus computational complexity) will be lower in smoother versions, but the loss of accuracy with regard to biomedical predictions remains to be investigated.
Table 1: Geometrical details of the realistic artery (Level 1) considered in the study
Dimensions Symbol Value
Artery length $L$ 17.2 cm
Inlet cross-section area $A_{\text{in}}$ 2.3 cm2
Outlet cross-section area $A_{\text{out}}$ 1.2 cm2
2.2 Governing equations
Patient-specific and simplified models of the arteries are used to study the hemodynamics of blood flow under pulsatile flow conditions. In this respect, the incompressible form of the continuity and momentum equations are solved with realistic boundary conditions with appropriate viscosity models. The incompressibility constraint in the form of the continuity equation and the applicable momentum equations for a variable viscosity fluid are given below.
$\nabla\cdot\textbf{u}=0$ (1)
$\rho\left(\frac{\partial\textbf{u}}{\partial t}+\textbf{u}\cdot\nabla\textbf{u}\right)=-\nabla p+\nabla\cdot\bar{\bar{\tau}}$ (2)
Here u is the velocity vector, $p$ is pressure, and $\bar{\bar{\tau}}$ is the deviatoric part of the total stress tensor. The stress tensor is a function of the strain rate given as $\bar{\bar{\tau}}=\mu(\dot{\gamma})\bar{\bar{\gamma}}$ , where $\bar{\bar{\gamma}}=\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$. The magnitude of the strain rate is given as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}:\bar{\bar{\gamma}}}$. Here $\mu$ is the apparent viscosity of blood based on the selected viscosity model.
2.2.1 Boundary Conditions
To solve the governing equations for the descending aorta and its simplified geometric versions, the following boundary conditions are applied:
Figure 2: Pulsatile velocity waveform incorporated at the inlet of the artery along with the pulsatile pressure waveform at the outlet of the descending aorta. [25] .
  1. 1.
    Inlet conditions: A uniform inflow velocity profile is imposed at the inlets of all geometries, combined with a pulsatile waveform to replicate the periodic nature of blood flow under physiological conditions. This pulsatile waveform captures the transient characteristics of flow and ensures a realistic simulation of the hemodynamic environment. The pulsatile velocity profile is derived from physiological measurements reported by Mills et al [25]. The derived profile is fitted by a Fourier series of up to 18 terms to apply the inlet boundary condition. To fully allow the inflow to develop into a Womersley profile, the inlet region of the artery is extended by a length of $10D$, where $D$ represents the characteristic diameter of the artery. This extension minimizes upstream flow irregularities owing to flow development and ensures a realistic inflow condition in the computational domain of interest. Similarly, an extension of $5D$ is added to the outlet to mitigate numerical reflections of an outflow boundary condition and maintain a smooth flow transition at the exit plane.
  2. 2.
    Outlet conditions: Realistic pulsatile gauge pressure is applied at the outlets of the geometries, as depicted in Figure 2. This boundary condition mirrors the dynamic pressure variations observed in the cardiovascular system, which are critical for accurately capturing the pressure-driven nature of blood flow. The pulsatile pressure profile is derived from physiological measurements carried out by Mills et al [25], ensuring that they impose real-world conditions. These outlet conditions play a key role in maintaining the accuracy of flow characteristics, especially in scenarios where downstream vascular resistance and compliance significantly affect the hemodynamics.
  3. 3.
    Wall conditions: The arterial walls in all geometries are assumed to be rigid, i.e. they do not deform under the influence of pressure and shear loading. A no-slip boundary condition is enforced at the walls, requiring the velocity at the wall surface is zero. This assumption is widely used in computational blood flow simulations.
2.2.2 Blood viscosity models
The selection of the viscosity model is essential to simulate realistic blood flow characteristics. Blood is a non-Newtonian fluid, but since shear rates are smaller in vessels of large diameters, non-Newtonian effects are less pronounced here. For such arteries, earlier studies have shown only minor differences in predictions from Newtonian and non-Newtonian models in tubular geometries. For the geometry considered in the present study, surface perturbations are present in Level 1 and differences in predictions across models may be expected. Thus, in this study, we consider the Newtonian and two non-Newtonian models for blood viscosity. The selected viscosity models are discussed below.
  1. 1.
    Newtonian
    The Newtonian model [26] is widely employed in blood flow simulations, primarily due to its linearity and the resulting computational efficiency. It is expected that for arteries of relatively larger size, Newtonian model predictions are negligibly different from the non-Newtonian ones. In the present study, blood viscosity is taken to be a constant, namely
    $\mu_{app}=0.0035\;\text{Pa s}$ (3)
    It is appropriate for a body temperature of 310 K.
  2. 2.
    Carreau-Yasuda
    The Carreau-Yasuda [27] is one of the most widely used Generalized Newtonian fluid (GNF) model. For viscoelastic fluids, the Carreau-Yasuda model is more realistic than the Power-law model since it shows a plateau at both low and high shear rates. Specifically, blood demonstrates near-constant viscosity at high shear rates, consistent with the assumptions of a Newtonian fluid under such conditions. The Carreau-Yasuda viscosity model is given as:
    $\mu_{app}=\mu_{0}+(\mu_{0}-\mu_{\infty})(1+(\lambda\dot{\gamma})^{P})^{\frac{n-1}{P}}$ (4)
    In this context, $\mu_{0}$ and $\mu_{\infty}$ are the saturation values of viscosity at low and high shear rates, respectively. The parameter $P$ is known as the Yasuda exponent, with both $P$ and $n$ determining the degree of nonlinearity. The time constant $\lambda$ regulates the range of strain rates at which the transition to a power-law type regime is obtained. In a way, $\lambda$ is related to the relaxation time of the fluid. The model simplifies to the Newtonian fluid when the time constant is zero ($\lambda$ = 0). Table 2 lists the parameters used for the CarreauYasuda model for blood in this study.
    Table 2: Parameters considered for the Carreau-Yasuda blood viscosity model at a body temperature of 310 K
    Carreau-Yasuda Viscosity Parameters Value
    $\mu_{0}$ 0.056 Pa·s
    $\mu_{\infty}$ 0.0035 Pa·s
    $\lambda$ 3.313 s
    $n$ 0.3568
    $P$ 2
  3. 3.
    Apostolides and Beris
    The recently developed viscosity model by Apostolides and Beris [28] is based on the Casson viscosity model. The parameters are functions of the constituents of blood like hematocrit and fibrinogen, whose values are adjusted using various flow-related data available in the literature. The apparent fluid viscosity is given by the following equation:
    $\mu_{app}=\left(\sqrt{\frac{\tau_{o}}{\dot{\gamma}}}+\sqrt{\mu_{c}}\right)^{2}$ (5)
    Here, $\dot{\gamma}$ is the magnitude of the strain rate while the yield stress $\tau_{0}$ is given as follows:
    $\tau_{0}=\begin{cases}(\phi\ -\phi_{c})^{2}[0.508c_{f}+0.4517]^{2},&\text{if $\phi>\phi_{c}$};\\ 0,&\text{if $\phi\leq\phi_{c}$};\end{cases}$ (6)
    A threshold value of hematocrit is required to calculate the yield stress; this threshold is a quadratic function of the fibrinogen concentration $c_{f}$.
    $\phi_{c}=\begin{cases}0.312c_{f}^{2}-0.468c_{f}+0.1764,&\text{ $c_{f}<0.75$};\\ 0.0012;&\text{ $c_{f}\geq 0.75$};\end{cases}$ (7)
    Figure 3: Blood viscosity models selected for simulating the blood flow in the descending aorta. Newtonian, Carreau Yasuda, and Apostolides and Beris models are shown.
    The Casson viscosity, i.e., $\mu_{c}$, is also obtained by using the experimental data available in the literature and follows an Arrhenius dependence on temperature, given as:
    $\mu_{c}=\mu_{p}(1+2.0703\phi+3.722\phi^{2})\exp\left[-7.0276\left(1-\frac{T_{0}}{T}\right)\right]$ (8)
    At fully developed conditions, the local shear rate is zero at the center of a smooth circular channel. This local shear rate will make the apparent viscosity infinitely large, leading to an instability in the numerical solver. To ensure the stability of the solver at fully developed conditions, Miller and Morris [29] suggested that the shear rate magnitude is regularised as $\dot{\gamma}=\dot{\gamma}+\dot{\gamma}_{r}$, where the non-local shear rate $\dot{\gamma}_{r}<<\dot{\gamma}$ and given as
    $\dot{\gamma}_{r}=\frac{\epsilon U_{max}}{L_{c}}$
    Here, $U_{max}$ is the maximum velocity attained in the Newtonian viscosity model, $\epsilon=0.01$ is the regularization parameter, and $L_{c}$ is the characteristic length. In the present study, it is chosen as the inlet diameter of the Level 1 artery. The parameters selected are appropriate for a normal human body temperature of 310 K and summarized in Table 3.
    Table 3: Parameters considered for the Apostolidis and Beris blood viscosity model
    Blood Parameters for Apostolidis and Beris Viscosity Model Value
    Average hematocrit concentration, $\phi$ 0.4
    Temperature of Blood Sample, $T$ 310 K
    Reference Temperature, $T_{0}$ 296.1 K
    Plasma viscosity, $\mu_{c}$ $1.67\times 10^{-3}$ N·s/m2
    Density of blood, $\rho$ 1060 kg/m3
    Fibrinogen concentration, $c_{f}$ 0.125 g/dl
3 Flow distribution in a descending aorta
3.1 Grid Independence Test (GIT) and Validation
Simulations of the flow distribution in the descending aorta in patient-specific (Level 1) and its simplified versions (Levels 2 and 3) are performed using the finite volume method within Ansys-Fluent, a commercial flow solver. For the grid independence test for the three geometries, the steady flow problem is solved on four refined meshes (M1, M2, M3, and M4). The solver used in this study is also validated with the results of Nagargoje et al. [30] for a symmetric bifurcation with pulsatile inlet boundary conditions. The pulsatile waveform and the geometry of interest are shown in Figure 4. In Figure 5(a), the average velocity obtained is shown as a function of the inverse of the mesh element size. Following grid independence, the M3 mesh is selected for all further simulations. Validation against the work of [30]is shown in Figure 5(b). The $z$-component velocity is plotted on the centerline near the bifurcation for three-time instants of the oscillatory inflow condition. The validation results are seen to be satisfactory, which provides a check on the correctness of the solver.
Figure 4: (a) Oscillatory velocity inlet profile. (b) symmetric bifurcation geometry used by Nagargoje et al. [30]
Next, flow metrics and hemodynamic parameters are analyzed to understand the implications of the geometric simplifications for the three viscosity models. Quantities of interest are the velocity profiles and wall shear stress, along with the hemodynamic parameters such as the time-averaged wall shear stress (TAWSS) and Oscillatory Shear Index (OSI). Analysis in terms of the Q-criterion for strain rate and rotation are also reported.
Figure 5: (a) Mesh independence test is performed for Levels 1-3 of the descending aorta. The average velocity at the outlet is presented as a function of the inverse of the mesh element size. Here, $e$ is the mesh element size, and a higher inverse element size indicates a finer mesh. (b) The $z$-velocity component (w) is plotted at time instants $t_{1},t_{2},$ and $t_{3}$ at the bifurcation of the geometry and matched against comparable results of Nagargoje et al. [30]
3.2 Velocity
This section presents the cross-sectional velocity distribution during the systolic and diastolic phases over various planes of the descending aorta for the three viscosity models and Levels 1-3 geometries. The velocity data are shown in Figures 6 and 7 for the 4th cycle of simulation since the flow was shown to exhibit clear periodicity after the third cycle, indicating a fully developed dynamic steady state. Specific time instants from the cycle, as highlighted in left-most columns of Figures 6 and 7 are used for discussion.
The systolic phase of the waveform (Figure 2) includes three key instants: $t_{1}$ = 0.2 s in the initial instant of increase in inflow velocity, $t_{2}$ = 0.3 s is the peak, and $t_{3}$= 0.45 s is the deceleration phase. The diastolic phase is analyzed at two time instants, $t_{4}$ = 0.6s and $t_{5}$ = 0.7s, representing flow fluctuation towards the end of the cycle. This section further describes the velocity characteristics over the cross-section in the form of centerline plots along with streamlines, contours, and vectors for the selected viscosity models and geometric simplification levels. In Figures 6 and 7, the distribution of the axial velocity component (denoted by $u$) is shown at the centerline of the cross-section located at a distance of $z=16D$ from the extended inlet as a function of the normalized $y$-coordinate. We show a comparison across two factors; in the first set of figures (Figure 6), we focus on comparing the viscosity model for a particular level of simplification, which shows subtleties in model predictions. In Figure7, we show a comparison across the level of simplification for any given viscosity model. This completes the picture and shows how modeling aspects are coupled with geometry details. In Figure 6, each level of geometry simplification is shown in a column, while rows represent distinct phases of the pulsatile flow cycle. The velocity profiles, normalized across the channel width, are depicted by red, blue, and black lines for the Newtonian, Carreau-Yasuda, and Apostolidis-Beris models, respectively. The axial velocity component distribution is shown at the selected time instants for the individual viscosity models for a particular geometry simplification level. An asymmetric velocity distribution is observed for Level 1 only. This is expected since the perturbation and undulations across the circumference at any axial position are retained in Level 1, which leads to an asymmetric flow in succeeding cross-sections. Interestingly, in Level 1, the flow profiles are nearly identical for all phases of the cycle for the Newtonian and Carreau-Yasuda viscosity models. This similarity is expected since the Newtonian and Carreau-Yasuda models predict similar viscosities at high shear rates (Figure 3). The velocity predictions for the Apostolidis and Beris viscosity model are slightly larger than the other two for all the phases. For Level 2 (Figure 6(b)), the geometry is simplified and made symmetric around the circumference, and symmetric flow profiles are observed for all the viscosity models at every phase of the pulsatile flow. Interestingly, at this level, the two non-Newtonian models - Carreau-Yasuda and Apostolidis and Beris- predict similar velocity profiles, whereas the Newtonian predictions show an overshoot near the centerline. Note that such differences at the centerline are known between Newtonian and non-Newtonian models for flows in a smooth duct where the latter needs a resolution of the shear rate singularity within the definition of apparent viscosity. For the simplest Level 3 in Figure 6(c) which is a tapered smooth tube, the predictions across viscosity models are largely symmetric and similar to each other, owing to the symmetry of the geometry.
Next, in Figure 7, we show the velocity comparison across geometry simplifications for each viscosity model. The data shows an accuracy check for such simplifications across different viscosity models considered in the present study. Columns (a), (b), and (c) depict the velocity profiles for the Newtonian, Carreau-Yasuda, and Apostolidis and Beris viscosity models, respectively. As mentioned earlier, the rows represent distinct phases of the pulsatile flow cycle. Largely, the trends agree with each other, except for the asymmetry that is only visible for the patient-specific geometry (Level 1). The largest differences are observed at the start of the cycle (first row) and during the deceleration phase (third row). Interestingly, for the deceleration phase, only the predictions with the combination of Level 3 and Newtonian fluid are significantly more significant than the others. Thus, with respect to velocity profiles, the simplified geometry can capture the trends quite well for all viscosity models, except for the expected asymmetry in Level 1.
In what follows, further analyses in terms of streamlines, velocity contours, and vectors are provided. Figure 8a shows velocity magnitude contours at various cross-sectional planes in the descending aorta in the Level 1 geometry for various phases of the inflow waveform. Additionally, Figure 8b displays the velocity magnitude combined with streamline plots over the medial plane of the Level 1 aorta. These details are presented here for the Apostolidis and Beris viscosity model. The key trends remain similar to the other two models and hence, are not repeated. During the systolic phase, the inflow increases steadily, reaching a peak corresponding to the maximum flow rate. This progressive increase in inflow elevates wall forces, consequently enhancing wall loading. In contrast, the diastolic phase exhibits a much more complex flow behavior, driven by the adverse pressure gradient that develops as the inflow decelerates in time. The adverse gradient generates regions of backflow where the flow reverses direction near the walls.
Figure 6: Axial velocity component shown for the three different viscosity models in (a) Level 1, (b) Level 2, and (c) Level 3 geometries.
Figure 7: Axial velocity component shown for the original and two simplified geometries (Levels 1-3) for (a) Newtonian, (b) Carreau-Yasuda, and (c) Apostolidis and Beris viscosity models.
These regions are particularly pronounced in areas with irregular geometry or abrupt changes in cross-section and surface perturbations. This backflow disrupts the unidirectional flow and interacts with the surrounding fluid, leading to flow separation and the formation of recirculation zones. The interplay between backflow, flow circulation, and pressure gradients indicates the sensitivity of the flow field to boundary conditions and geometric variations and reinforces the importance of retaining details in the geometry. Thus, a prediction accuracy loss is expected with geometry simplifications, some of which are examined next in Figure 9.
Streamlines in the three geometries at all time instants are shown in Figure 9 with the Apostolidis and Beris viscosity model. It is observed that during the systolic phase, all the streamlines are unidirectional and no flow circulation is observed. However, as the flow enters the diastolic regime, flow circulations appear in Level 1 as well as Level 2 but are completely absent in Level 3. These flow recirculation zones are a result of the interplay of the geometrical perturbations and adverse pressure gradient conditions. With the absence of perturbations in Level 3, such zones of backflow are absent. Reversed flow zones are also larger in Level 1, relative to Level 2, owing to geometric variations coupled with a loss of symmetry. Thus, as the artery geometry is simplified, the intensity of circulation decreases and disappears in the limit of the smooth geometry approximation, namely, Level 3. Flow circulation patterns are important from a medical point of view since they affect the wall shear stress (WSS) and pressure loading. The circulation region being stagnant will promote particulate deposition leading to thrombosis. Hence, delineating such regions is important from a biomedical perspective.
a
b
Figure 8: (a) Velocity magnitude contours at various time instants for Level 1 artery and (b) streamlines superimposed on the velocity magnitude contour over the median plane for the Apostolidis and Beris viscosity model. (i)-(iv) show distinct phases of the vascular waveform.
Figure 9: Velocity contours superimposed with streamlines are shown at selected time instants for the Apostolidis and Beris viscosity model for the three geometry levels 1-3. Within (i)-(v), Levels 1 to 3 are shown from left to right.
3.3 Hemodynamic Parameters
3.3.1 Wall shear stress (WSS)
Wall shear stress (WSS) plays a crucial role in regulating endothelial function and has been widely recognized as a key factor in the development of atherosclerosis. Endothelial cells sense the wall shear stress via multiple responsive cell components, and activation (or) deactivation of intra-cellular pathways leads to gene and protein expression regulation[31]. This continuous process is important for the remodeling, development, and growth of the vascular walls. Moreover, clinical significance of wall shear stress lies in its ability to predict the risk of future cardiovascular events, making it an important biomarker in cardiovascular health. Mathematically, the WSS can be calculated with the help of velocity gradients. $(\nabla\textbf{u})$ as follows:
$\bar{\bar{\tau}}_{w}=2\mu_{app}\bar{\dot{\gamma}}$ (9)
The strain rate ($\bar{\bar{\gamma}}$) is given by the sum of the velocity gradient and the transpose of the velocity gradient $\frac{1}{2}(\nabla\textbf{u}+\nabla\textbf{u}^{T})$ at the wall. The magnitude of the strain rate tensor field can be calculated as $\dot{\gamma}=\sqrt{2\bar{\bar{\gamma}}\colon\bar{\bar{\gamma}}}$. The wall shear stress is calculated for the patient-specific and simplified geometries (all levels) and presented in the form of contours.
Figure 10: Wall shear stress (WSS) contours superimposed with streamlines are shown at the selected time instants of the vascular cycle for the Apostolidis and Beris viscosity model. Within (i) to (v), Levels 1 to Level 3 are shown from the left.
The wall shear stress (WSS) contours are shown in Figure 10 for all the time instants of the pulsatile cycle for the original and other geometric simplification levels while adopting the Apostolidis and Beris viscosity model. The features of WSS in Level 1 are particularly significant in predicting the onset and progression of disease, as they provide detailed insights, as well as the identification of the critical regions. Figure 10 shows how the simulation accuracy deteriorates by simplification of the geometry. Firstly, for Level 1, with all perturbations preserved, we observe high variability in WSS at the peak instant of flow. At all times, the surface perturbations and the asymmetry generate several high (and low) stress zones of WSS. Some of these variabilities are captured in Level 2, although the asymmetry of Level 1 is lost. The high and low-stress zones appear as bands around the circumference, with an overall reduction of the variations observed in the WSS distributions of Level 1. This is most pronounced at the time instant of peak velocity. However, at Level 3 (tube of decreasing area), the WSS profile is smooth and all the variability is lost. Thus, Level 2 might still provide satisfactory predictions of WSS, which is entirely lost in Level 3.
3.3.2 Time averaged wall shear stress (TAWSS)
Along with WSS, time-averaged wall shear stress (TAWSS) is a primary risk parameter to monitor the development and progression of atherosclerosis, which can further lead to various types of cardiovascular diseases. Both low and high-wall shear stress promotes pro-inflammatory and anti-inflammatory responses by the endothelial cells [31]. An average WSS over the cardiac cycle is the time-averaged wall shear stress (TAWSS), which further helps identify the high and low WSS zones. Mathematically, TAWSS is defined by the following relation:
[…]
Here, $T$ is the time period of the cardiac cycle. Time-averaged wall shear stress (TAWSS) calculated for the fourth cycle of the simulations is shown in Figure 11 for the three levels of the geometry of the descending aorta and selected viscosity models. In Figure 11 (i),(ii) and (iii) TAWSS contours are plotted for Apostolidis and Beris, Carreau Yasuda, and Newtonian viscosity models, respectively. The trends are qualitatively similar for WSS discussed earlier. In Level 1, the perturbations and asymmetry of the geometry generate a significant amount of zones for high and low TAWSS. Some of these zones are retained in Level 2 with less asymmetry and a reduction in the shear stress magnitudes. A smooth variation of TAWSS is obtained in Level 3 which is a smooth geometry. The trends are similar across the three viscosity models. However, the perturbations in TAWSS are slightly different with the Apostolidis and Beris model, whereas the profiles are similar for the Newtonian and Carreau-Yasuda models. The values are slightly higher with the Apostolidis and Beris model in Level 1. Thus, again, concerning spots of extreme stress, Level 2 captures the essence partially while Level 3 shows a smooth variation.
Figure 11: Time average wall shear stress (TAWSS) shown for the three viscosity models for Levels 1,2 and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
3.3.3 Oscillatory shear index (OSI)
As discussed earlier, WSS levels activate and deactivate the endothelial cell response. In pulsatile flow, the shear stress near the artery wall fluctuates, making the endothelial cell deliver a pro-inflammatory response. The Oscillatory Shear Index (OSI) introduced by Ku et al. [32] provides a measure of these fluctuations. The OSI can be calculated as follows:
[…]
From the definition of OSI, the high-value region shows a high time-periodic oscillation of the wall shear stress. OSI varies between 0 and 0,5, the latter representing a purely oscillatory response.
Figure 12: Oscillatory shear index (OSI) shown for the three viscosity models for Levels 1,2 and 3 geometry of the descending aorta. (a) Apostolidis and Beris, (b) Carreau-Yasuda and (c) Newtonian viscosity model.
Figure 12b plots Oscillatory Shear Index Contours for the selected viscosity models. OSI contours depict the zone of shear oscillation and its intensity. In Figure 12, the OSI distributed zones in the Level 1 artery can be seen with hotspots of high and low OSI values, signifying the nature of shear stress oscillations. The region with high OSI values is more prone to a pro-inflammatory response by the endothelial cells, whereas the low value of the OSI zone indicates unidirectional shear stress; therefore, an anti-inflammatory response for the cell is obtained. As the complexity of the descending aorta gradually decreases, the OSI zones are also averaged uniformly as the geometrical features are toned down in the subsequent levels of simplification. Lastly, a uniform OSI distribution is observed in Level 3, as no geometrical disturbances are present. It is clear that the asymmetry of Level 1 is lost in both Levels 2 and 3 due to the geometric simplification. Thus, a fair comparison of the time-averaged hemodynamic indicators across geometric levels is shown in Fig. 13 for TAWSS (Fig. 13 (a)) and OSI (Fig.13 (b)) averaged over the circumference at each axial location. As might be expected, the predictions in Level 3 are a monotonic function, while predictions in Levels 1 and 2 show oscillations due to axial variations in the geometry. For all data shown, the Level 3 predictions represent an average behavior of the oscillatory predictions obtained for Levels 1 and 2. The usefulness of Level 2 is readily visible in these results since, mostly, it closely follows the predictions of Level 1. In the present discussion, the Level 1 predictions are averaged over the circumference and the Level 2 geometry is built by averaging diameters at various axial locations. The consistency of Levels 1 and 2 is visible even more for the non-Newtonian models and shows some differences for the Newtonian fluid, especially for OSI. These results bolster the proposition that a somewhat smoother geometry (like Level 2) will be able to capture the fluctuations in hemodynamic parameters, although the local asymmetry will be lost. As discussed later in Section 3.5, this presents a valid ground for a speed-accuracy trade-off since convergence in the smoother geometries is faster and yet captures the essence of the predictions of Level 1. However, only Level 1 (patient-specific geometry) can be used without any further simplifications in applications where the details of local asymmetry are required.
Figure 13: Comparison of TAWSS and OSI for (i) Newtonian, (ii) Carreau-Yasuda, and (iii) Apostolidis and Beris models. (a) TAWSS: Time-Averaged Wall Shear Stress. (b) OSI: Oscillatory Shear Index for the geometric simplification levels 1-3 of the descending aorta. Results of Level 1 shown here are averaged over the circumference for each axial location.
For TAWSS, Level 2 is able to provide predictions close to those of the circumferential averaged Level 1 data. However, Level 3 is only able to show a monotonic function representing an overall average trend. It lacks the oscillation and fluctuations that are caused due to area fluctuation, characteristic of Levels 1 and 2. It is possible to ask, from the prediction of TAWSS of Level 3 and knowing the area changes, is it possible to recover the prediction of Levels 1 and 2? The following approach is adopted. Shear stress is given by
$\text{TAWSS}\sim\mu_{app}\frac{\partial u}{\partial z}$
for flow in a duct. In terms of the order of magnitudes, the expression can be written as
$\text{TAWSS}\sim\mu_{app}\frac{U}{D}\frac{\partial u*}{\partial z*}$
Here, symbols $U$ and $D$ are the characteristic velocity and diameter at any location within the duct. The average velocity at any location can be estimated by knowing the volumetric flow rate and local cross-sectional area. Since we know the total volumetric flow rate and area at any location, $U$ and $D$ can be estimated for local perturbations while $\frac{u*}{z*}$ can be assumed to be unaffected. At a high shear rate, $\mu_{app}$ can be assumed to be a limiting value (Figure 3). Thus, TAWSS can be estimated locally by estimating the characteristic velocity and diameter at every cross-section. These estimates are shown in Figure 13(a) for all viscosity models by black dashed lines.
These results clearly indicate the success of the estimated TAWSS across all models, with non-Newtonian models performing slightly better. Thus, knowing the results of the simulation for Level 3 and the prescribed area perturbations, the realistic local TAWSS in a patient-specific geometry can be readily estimated.
3.4 $Q$ criterion
Another critical parameter for studying strain rate and rotational characteristics is the variation of the $Q$ criterion. $Q$ criterion is based on the velocity gradient tensor, which is given as the summation of a symmetric and anti-symmetric part as follows:
$\nabla\textbf{u}=\frac{\nabla\textbf{u}+\nabla\textbf{u}^{T}}{2}+\frac{\nabla\textbf{u}-\nabla\textbf{u}^{T}}{2}$
The first term is the symmetric part and represents the rate of strain tensor (${\bar{\bar{\gamma}}}$). The second term represents the anti-symmetric part of the vorticity tensor (${\bar{\bar{\Omega}}}$). The $Q$ criterion is then calculated as:
$Q=\frac{||\bar{\bar{\Omega}}||-||\bar{\bar{\gamma}}||}{2}$ (12)
A positive $Q$ value shows the region with a dominance of rotational flow. Similarly, a negative value denotes a zone where rotational flows are not dominant. In Figure 14a, an iso-surface of the $Q$-criterion is shown for $Q$ = 0.1. Similarly, in Figure 14b, it is shown for $Q$ = -0.1. It is clear that the simplification of the artery leads to a uniform flow distribution. In Level 1, the $Q$ criterion shows a non-uniform asymmetric distribution; as the geometry is simplified to Level 2, the distribution becomes symmetric and breaks into zones. In Level 3, only a smooth distribution is obtained. In this discussion, the value of the $Q$ function is selected in such a way as to fill the descending aorta. It is observed that in the systole phase, the vortical structure for Level 1 is distributed in all regions. However, for Level 2, the same structure breaks and spreads towards the periphery of the wall. For Level 3, the vortical structures are uniformly distributed over the entire flow space. As the flow rate increases, the structure becomes stronger. However, In the diastolic phase, the strain rate grows and dominates across all levels due to a lower flow rate. Further, separation and breakage of the flow structures are also observed. A similar trend is also seen for the contours of $Q$ = -0.1, as shown in Figure 14b. The nature of the distribution of the $Q$ function is associated with wall shear stress and, correlates closely with those of the hemodynamic indicators, namely, TAWSS and OSI.
a
b
Figure 14: Iso-surfaces of the $Q$ criterion are shown for dimensional values of (a) 0.1 and (b) - 0.1 values for the three geometries simulated with the Apostolidis and Beris viscosity model at the selected time instants of the vascular waveform. Within time instants of (i) - (v), the three levels are arranged from left to right.
3.5 Speed Accuracy trade-off
In the preceding sections, we have shown details of the accuracy aspects arising from various simplification levels. As discussed, Level 2 mostly captures the local effects for all parameters, but the local asymmetry visible in the original Level 1 is lost. Level 3 is unable to capture any of the local effects and only shows a “representative average” of results obtained in Level 1. However, for the wall shear stress, simulation data of Level 3 can be reconstructed using scaling principles, which agree well with the detailed calculations of Level 1. Thus, each level has its own utility and is able to provide limited but useful information about the flow field. In the present section, we discuss the reduction in computational time with geometric simplification. In Level 2 geometry, we used 25 individual slices over the length of the artery, each slice being a cylindrical patch. To perform an analysis of the effect of this number, we reconstructed the entire artery by extracting the cross-sectional area of the Level 1 geometry along its length. We then developed multiple equivalent geometries consisting of 25, 20, 15, 10, and 5 slices apart from Level 3, which does not have any slices. Using a grid-independence test, we analysed all reconstructed geometries with Newtonian blood viscosity under steady flow conditions, employing a uniform inlet velocity.
In Figure 15(a), we show the friction factor for all reconstructed geometries, including the Level 1 geometry, as a function of the mesh element size. For Level 1 geometry, the converged mesh size appears to be around 0.6. Similarly, the friction factor for all other simplified geometries is plotted, with convergence occurring at about an element size of 0.8. The results indicate that the simplified geometries require coarser meshes to achieve grid convergence compared to the Level 1 geometry. This finding aligns with expectations, as the simplified geometries are symmetrical along their axis and less complex, making convergence attainable on coarser meshes. Consequently, the simplified arteries are expected to yield faster simulations than Level 1. Apart from the asymmetry, most of the local flow aspects are resolved by the Level 2 geometry. Thus, we can perform simulations to a similar level of accuracy using a coarser mesh, thereby increasing the computational efficiency of the simulation.
Figure 15: Speed-accuracy analysis is carried out by performing (a) grid independence test for simplified geometries up to 25 slices and comparing their respective friction factors. (b) The maximum number of iterations required to reach the tolerance of 1e-5 in velocity is plotted as a function of the element size. (c) Maximum number of iterations is plotted for the Apostolidis and Beris and Newtonian viscosity models for Level 2 (with 25 slices) and Level 3 simplification.
Further, to show the speed increase with geometric simplification, we show the maximum number of iterations required to reach a prescribed convergence criterion as a function of the element size, in Figure 15(b). The data reveal that simplified geometries with 5 to 20 slices require approximately 150 iterations. In contrast, the simplified artery with 25 slices and the geometry with zero slices need around 300 iterations for the smallest element size and 200 iterations for the largest element size. For Level 1 geometry, the number of iterations increases as the element size is decreased, which is consistent with expectations for unstructured meshes over complex three-dimensional geometries. Thus, there is a clear trend of improvement in terms of the number of iterations (and thus, the speed of computations) as we move towards simplified geometric representations. The difference is as large as four times between Level 1 and Level 2 (with 20 or fewer slices) at their optimal element sizes.
In Figure 15(c), the Apostolidis and Beris viscosity model is compared with the Newtonian viscosity model with respect to the maximum number of iterations as a function of element size for Level 2 and Level 3 geometries. A significant difference is observed when the viscosity model becomes non-Newtonian. For the converged mesh, nearly 5000 iterations are required for the non-Newtonian viscosity model as compared to the Newtonian. Overall, geometric simplification reduces the number of iterations required, nonlinearity of the viscosity model increases the number of iterations, while grid refinement increases the number of iterations, the discussion being carried out on meshes with acceptable grid quality.
4 Discussion
The study provides two broad observations. First, geometric simplifications can be useful for generating fast but approximate insights, especially when time constraints exist in clinical decision-making. The approximate solutions can often be improved using suitable scale analysis. However, for a detailed analysis, retaining complete geometrical fidelity is essential. Secondly, for larger arteries, the Newtonian model can provide qualitatively similar predictions faster than the non-Newtonian ones in terms of CPU time. The findings also emphasise the importance of combining accurate hemodynamic indicators such as time-averaged wall shear stress and the $Q$-function with newly developed blood viscosity models for better prediction of blood-tissue interaction. These insights could aid in designing personalised treatments, evaluating surgical interventions, and improving diagnostic accuracy in detecting the onset of cardiovascular diseases.
5 Conclusions
A detailed study on a patient-specific geometry of the descending aorta with two simplified representations and three viscosity models is conducted. The geometry is extracted from CT scan data shared by a hospital and reconstructed using image processing tools. Realistic time-dependent boundary conditions are applied at the inlet and outlet in terms of pulsatile velocity and outlet pressure, respectively. For the geometrical approximation, the original artery (termed Level 1) is further simplified into two more levels, denoted as Level 2 and Level 3. Level 2 is based on the cross-sections of Level 1, which is divided into 25 sections, with each part having a representative average diameter. Thus, Level 2 captures some of the undulations and area fluctuations of Level 1 but loses the local asymmetry of the original. Level 3 represents a smooth duct based on the inlet and outlet cross-sectional area of the actual artery. The calculations are performed with three blood rheology models – Newtonian, Carreau-Yasuda, and Apostolidis and Beris. Various aspects of blood flow for all the levels of the descending aorta geometry (with selected blood viscosity models) are studied. The following conclusions are arrived at in the present work.
  1. 1.
    As the geometrical complexity is averaged out, the overall flow field becomes uniform, with a reduction in the strength of the recirculation zones near wall perturbations. For Level 1 geometry, the flow field shows a significant number of recirculation zones near the walls. Some of these zones are visible in Level 2 as well, though with reduced intensity. No such zone of reversed flow is visible in Level 3, where the flow field is practically unidirectional.
  2. 2.
    The simplification of the geometry affects the strain rate distribution, further influencing the wall shear stress (WSS). Simulations show that Level 2 geometry could approximate the WSS distribution reasonably well, capturing major patterns and magnitudes while missing the local asymmetry in the stress profiles. However, Level 3 shows only a smooth variation of the WSS without showing undulations.
  3. 3.
    The variations of the hemodynamic indicators such as TAWSS and OSI bolster these earlier assertions further. Level 2 predictions are particularly close to those obtained in Level 1, especially when both sets of data are averaged over the circumference. However, Level 3 predictions are without any fluctuations and are, at best, a representative average.
  4. 4.
    A simple estimate of the actual WSS in the original artery can be obtained by knowing the prediction from Level 3, scaled by the area fluctuation. The estimates thus retrieved are pretty close to the ones obtained with Level 1.
  5. 5.
    The speed test showed that computations with Levels 2 and 3 could be about 4 times faster than Level 1. The Newtonian model needs about 10 times fewer iterations than the Apostolidis and Beris model for a comparable convergence level and grid refinement.
Acknowledgment
Authors acknowledge the National Supercomputing Mission (NSM) for providing computing resources of ’PARAM Sanganak’ at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. Also, we would like to thank the computer centre (www.iitk.ac.in/cc) at IIT Kanpur for providing the resources to carry out the reported work.
Author Declaration
The authors have no conflicts to disclose.